# Deferred NAM: Low-latency Top-K Context Injection via Deferred Context Encoding for Non-Streaming ASR

### v1

@@ 1 Introduction
In this work, we optimize context encoding by splitting it into two passes. We make the simplifying assumption that our ASR system is non-causal, i.e., the entire audio; we argue that this choice is not overly restrictive, as two-pass ASR systems combining a causal ASR system to produce streaming results with a non-causal ASR system for producing the final result have become common, e.g., as in Narayanan et al. (2021). With the speaker’s entire audio available before context encoding begins, we can split context encoding into two phases. First, we encode all contextual phrases with an extremely lightweight encoder, and use the resulting encodings to determine the phrases that are most likely to occur in the audio. We embed only the most-likely phrases with a more powerful — but more expensive — encoder, and use the output of this “deferred” encoder for biasing.
@@ 2 Related Work
Attention-based biasing Pundak et al. (2018); Chang et al. (2021); Munkhdalai et al. (2022), in which the ASR network learns to use inference-time context through attention, and thus requires no separate contextualization components and can be optimized end-to-end by standard backpropagation, has become a popular method for contextualization of end-to-end ASR models. The core technique has spawned several distinct lines of complementary research. There are threads of work on improving data efficiency by adapter-style training Sathyendra et al. (2022) or improving training on synthetic audio derived from text-only data Naowarat et al. (2023). Work on precision improvements seeks to lower the rate of over-biasing through hierarchical or gated attention Han et al. (2022); Munkhdalai et al. (2023); Wu et al. (2023); Tong et al. (2023a); Xu et al. (2023a); Alexandridis et al. (2023) or through slot triggering Lu et al. (2023); Tong et al. (2023b); often, such techniques can improve not only quality but also per-step inference run time Munkhdalai et al. (2023); Wu et al. (2023); Yang et al. (2023). Notably, Tong et al Tong et al. (2023a) improve WER through an auxiliary slot-level cross-entropy loss. We do not use slots for selecting context to bias, as we have found it possible to get high performance with hierarchical attention alone; however, in this work, we do apply a phrase-level, rather than slot-level, cross-entropy loss during training to improve WER. Further we extend the technique to a novel wordpiece-level cross-entropy loss. Further work aims to improve quality on biased utterances by providing the context encoder with more information than just the graphemic biasing phrases, including phoneme-level features Bruguier et al. (2019); Hu et al. (2019); Chen et al. (2019); Pandey et al. (2023), and augmenting the context encoder with semantics-aware embeddings from a pretrained BERT model Fu et al. (2023). Attention-based biasing can also be complemented by shallow fusion with (contextualized) language models Xu et al. (2023b) or combinations of language models and contextualized rescoring Dingliwal et al. (2023) for further quality improvements.
@@ 5.3 Exploration: NO_BIAS filter
We investigated gating the WP attention with the phrase-level NO_BIAS token, where a bias phrase is deemed “active” if its per-frame logit is higher than that of the NO_BIAS token () at any frame. Method 1 (M1) filters inactive phrases before WP attention (equation 20). However, with fewer WPs to attend to, the remaining ones receive higher attention probabilities. This could degrade anti-context WERs if the accuracy of the NO_BIAS token is not sufficiently high, i.e., under condition , only 37.5% of utterances are deemed inactive for D3 at ANTI (3K). Therefore, we explored Method 2 (M2), which zeroes out the WP attention values of inactive phrases, leaving the probabilities of the WP keys unaffected (equation 21).

### v2

@@ 1 Introduction
In this work, we optimize context encoding by splitting it into two passes. We make the simplifying assumption that our ASR system is non-causal and can access the entire audio input; we argue that this choice is not overly restrictive, as two-pass ASR systems combining a causal ASR system to produce streaming results with a non-causal ASR system for producing the final result have become common, e.g., as in Narayanan et al. (2021). With the speaker’s entire audio available before context encoding begins, we can split context encoding into two phases. First, we encode all contextual phrases with an extremely lightweight encoder, and use the resulting encodings to determine the phrases that are most likely to occur in the audio. We embed only the most-likely phrases with a more powerful — but more expensive — encoder, and use the output of this “deferred” encoder for biasing.
@@ 2 Related Work
Attention-based biasing Pundak et al. (2018); Chang et al. (2021); Munkhdalai et al. (2022), in which the ASR network learns to use inference-time context through attention, and thus requires no separate contextualization components and can be optimized end-to-end by standard backpropagation, has become a popular method for contextualization of end-to-end ASR models. The core technique has spawned several distinct lines of complementary research. There are threads of work on improving data efficiency by adapter-style training Sathyendra et al. (2022) or improving training on synthetic audio derived from text-only data Naowarat et al. (2023). Work on precision improvements seeks to lower the rate of over-biasing through hierarchical or gated attention Han et al. (2022); Munkhdalai et al. (2023); Wu et al. (2023); Tong et al. (2023a); Xu et al. (2023a); Alexandridis et al. (2023) or through slot triggering Lu et al. (2023); Tong et al. (2023b); often, such techniques can improve not only quality but also per-step inference run time Munkhdalai et al. (2023); Yang et al. (2023). Notably, Tong et al Tong et al. (2023a) improve WER through an auxiliary slot-level cross-entropy loss. We do not use slots for selecting context to bias, as we have found it possible to get high performance with hierarchical attention alone; however, in this work, we do apply a phrase-level, rather than slot-level, cross-entropy loss during training to improve WER. Further we extend the technique to a novel wordpiece-level cross-entropy loss. Further work aims to improve quality on biased utterances by providing the context encoder with more information than just the graphemic biasing phrases, including phoneme-level features Bruguier et al. (2019); Hu et al. (2019); Chen et al. (2019); Pandey et al. (2023), and augmenting the context encoder with semantics-aware embeddings from a pretrained BERT model Fu et al. (2023). Attention-based biasing can also be complemented by shallow fusion with (contextualized) language models Xu et al. (2023b) or combinations of language models and contextualized rescoring Dingliwal et al. (2023) for further quality improvements.
@@ 5.3 Exploration: NO_BIAS filter
We investigated gating the WP attention with the phrase-level NO_BIAS token, where a bias phrase is deemed “active” if its per-frame logit is higher than that of the NO_BIAS token () at any frame. Method 1 (M1) filters inactive phrases before WP attention (equation 20). However, with fewer WPs to attend to, the remaining ones receive higher attention probabilities. This could degrade anti-context WERs if the accuracy of the NO_BIAS token is not sufficiently high, i.e., under condition , only 43.6% of utterances are deemed inactive for D3 at ANTI (3K). Therefore, we explored Method 2 (M2), which zeroes out the WP attention values of inactive phrases, leaving the probabilities of the WP keys unaffected (equation 21).